Find the slope-intercept from the equation of the line passing through the points (2,-4) and (4,6)
Answer by nerdybill(7384)

Find the slope-intercept from the equation of the line passing through the points (2,-4) and (4,6).
.
The "slope-intercept" form of the a line is:
y = mx + b
where
m is the slope
b is the y-intercept
.
Given two points of a line, we can calculate the slope using:
m = (y2-y1)/(x2-x1)
where
(x1,y1)=(2,-4) and (x2,y2)=(4,6)
.
m = (y2-y1)/(x2-x1)
m = (6-(-4))/(4-2)
m = (6+4)/(4-2)
m = (10)/(2)
m = 5
.
Now, use the "slope" (found above) AND one of the two points provided (say (2,-4)) to find 'b' (y-intercept). To do this, plug in what we know into:
y = mx + b
and solve for 'b'.
.
y = mx + b
-4 = 5(2) + b
-4 = 10 + b
-14 = b
.
Now that we found 'm' and 'b' plug it all back into:
y = mx + b
y = 5x + (-14)
y = 5x - 14 (this is what you're looking for)